OpenStudy (anonymous):

Which equations are true? Choose all answers that are correct. A. m∠2 + m∠1 = m∠3 B. m∠2 + m∠3 = m∠1 C. m∠3 – m∠2 = m∠1 D. m∠1 – m∠2 = m∠3

OpenStudy (anonymous):

can you help

OpenStudy (anonymous):

hello

rebeccaxhawaii (rebeccaxhawaii):

well 1 is made up of 2 angles what are they?

OpenStudy (anonymous):

?

rebeccaxhawaii (rebeccaxhawaii):

m1 is made up what two?

jhonyy9 (jhonyy9):

so than you check this posted image what you see there ? that angle 1 is the biger and that angle 1 is formed from angle 2 and angle 3 - yes ? so what mean these ?

OpenStudy (anonymous):

so a b are c are d

jhonyy9 (jhonyy9):

i think this will select it you after will understand this exercise - i will help you - ok. ?

OpenStudy (anonymous):

ok

jhonyy9 (jhonyy9):

so read please what i have wrote above and try solve it please

OpenStudy (anonymous):

Which equations are true? Choose all answers that are correct. A. m∠2 + m∠1 = m∠3 B. m∠2 + m∠3 = m∠1 C. m∠3 – m∠2 = m∠1 D. m∠1 – m∠2 = m∠3

jhonyy9 (jhonyy9):

do you understand these my posted words ? so like one example than 3 is formed from 1 and 2 so than 3 = 1+2 do you can using this example in this case of your above exercise please ?

OpenStudy (anonymous):

ok but whats the answer

jhonyy9 (jhonyy9):

you will give the answer sure after will understand this exercise read please my first answer and after this read please my second answer and with these knowladge will can solve it right sure easy

OpenStudy (anonymous):

Which equation can be used to find the value of a? A. a – 42 = 91 B. 42 + a = 91 C. 91 + a = 42 D. a – 91 = 42

jhonyy9 (jhonyy9):

your question was this with angles so first you need to solve it this - ok. ?

jhonyy9 (jhonyy9):

than you know that angle 1 is formed from angle 2 and angle 3 what mean this ? this is exactly like my above wrote example with 3 is formed from 1 and 2 and so 3 = 1+2 ok. ?